# Data-driven Forecasting of Deep Learning Performance on GPUs

### v1

@@ 3.3. Other Related Works
Performance modeling using analytical frameworks. Several works (Hong, 2009; Qi et al., 2017; Isaev et al., 2023) propose analytical models to predict GPU performance. Paleo (Qi et al., 2017) and Calculon (Isaev et al., 2023) decompose latency into computation and communication for distributed training. They estimate latency by measuring effective peak throughput with representative workloads on the target GPU, which requires access to the target GPU for estimation.
Cycle accurate simulators. Cycle-accurate simulators (Bakhoda et al., 2009) require detailed modeling of each GPU component. For every new GPU, the simulator must be updated to support new architectural features, which might not be readily available at launch. Proprietary architectural features can create a gap between real hardware and simulators. Additionally, cycle-accurate simulators are not suitable for end-to-end ML application execution due to their long simulation times. While certain works, such as sampled principal kernels (Avalos Baddouh et al., 2021), reduce simulation time, they still take at least hours to days to simulate modern DNN due to the need for modeling every detail of the hardware.
Machine Learning based approaches. Due to the large number of compute lanes in GPUs, they often observe linear performance trends for compute intensive kernels, and hence many prior works leverage regression models to predict the performance of a GPU architecture (Jia et al., 2012; Wu et al., 2015; Li et al., 2023). Some works require extracting low level hardware counter values to make predictions, however, these are unavailable unless the device is accessible (O’Neal et al., 2017). Such approaches cannot capture the complex intricacies of hardware features and software optimizations such as tiling for deep learning and do not support cross-generational performance prediction with GPUs due to the lack of hardware. Other works (Yang et al., 2022) propose graph neural networks to predict resource consumption (GPU utilization, memory utilization, and network TX/RX throughput) in distributed deep learning. However, this approach necessitates target GPUs for transfer training.
@@
4. NeuSight Forecasting for DNN Kernels
NeuSight forecasts the latency of executing a model on multiple GPUs in a server in three steps. The first and main step forecasts the latency of a single kernel that executes on the GPU atomically. The kernel leverages software and hardware optimizations for GPUs, namely tiling, parallel execution across SMs, on-chip caching, and HBM bandwidth. NeuSight uses architectural details such as memory size, memory bandwidth, number of SMs, L2 cache size, and peak FLOPS to predict the kernel’s latency through a machine learning approach. Our kernel estimation is inspired by how deep learning kernels are disassembled into multiple smaller tiles and mapped onto GPU hardware for execution. This approach contrasts with prior works that rely solely on compute or memory utilization to predict latency. Section 4.1 showcases the execution of a few representative DNN kernels on the GPU. NeuSight supports tensor operators such as General Matrix Multiplication (GEMM), element-wise operators such as addition and ReLU, and reduction-based operators such as Softmax and layer normalization.
In the second step, the per-operator latency is aggregated to predict the performance of the DNN workload based on its dataflow graph. In the third step, detailed in Section 5.1, the performance of per-GPU execution is combined with predicted network latencies to determine the performance of a DNN executing across multiple GPUs in a server.
@@ 4.1. Kernel Execution on GPUs
Figure 4. Each tile is distributed evenly across SMs and executed concurrently, in multiple number of waves. (NVIDIA, b)
Tiled execution. In this section, we focus on GEMM, a commonly found kernel in DNNs, but NeuSight can predict the latency of various operators to determine end-to-end performance. GEMM is a core building-block in deep learning layers, such as fully-connected, convolution, recurrent neural network, or self-attention found in modern large language models. Figure 3 shows a simplified view of a typical dataflow of GEMM on a GPU. Modern GPU libraries execute GEMM by first partitioning the output matrix into multiple identical tiles (Figure 3a). Each tile corresponds to a segment of the output matrix, loads the necessary input operands from input matrices, and computes the output elements. These tiles are mapped to each SM of the GPU and executed concurrently (Figure 3b). Typically the number of tiles which can be executed concurrently is limited by the number of SMs on the GPU, and the entire GPU kernel is executed in multiple waves of tile groups as depicted in Figure 4. Tiling strategy enables scalable execution of GEMM, by decomposing GEMM into multiple smaller workloads, which executes the identical computation on different input and output elements.
Fundamental performance laws of GPUs. Runtime of each kernel is bounded by a few features of a GPU, such as peak memory bandwidth or peak FLOPs. Roofline anlaysis (Williams et al., 2009) provides a straightforward method to estimate the approximate performance of a GPU kernel, taking into account the arithmetic intensity of the kernel. The intensity is computed by dividing number of flops ($flops_{k}$) by size of memory transactions ($mem_{k}$). The roofline bandwidth can be computed as follows:
(1) $\footnotesize K=\frac{flops_{k}}{mem_{k}};~{}~{}roofline_{BW}=min(K\times mem_{p},flops_{p})$
Where $flops_{p}$ is the peak FLOPs and $mem_{p}$ is the peak memory bandwidth of GPU. Roofline bandwidth represents the maximum achievable throughput of a kernel on a GPU.
@@
Latency hiding in GPU. Another key performance bound in GPU execution is the use of numerous threads to harness parallelism, achieving significantly higher throughput compared to traditional CPU architectures. GPUs can achieve the aforementioned peak hardware performance ($roofline_{BW}$) only when there are enough available software threads to effectively mitigate latency arising from data dependency or memory stalls by scheduling independent threads. Consequently, GPU performance relies heavily on the level of parallelism in the kernels, which is directly proportional to the number of threads that can be independently scheduled in each SM. Figure 5 illustrates the correlation between the number of waves available in each SM and the achievable throughput. As the number of waves increases, each SM is assigned more waves, resulting in a greater number of threads per SM. This allows for effective hiding of stall latency in each thread, thus leading to higher performance.
@@ 4.2. Kernel-wise Prediction
Tile-granularity prediction. For each DNN kernel we break it down into multiple identical tiles, instead of directly predicting the latency of the kernel. This insight is motivated from prior works (Yu et al., 2021), where a machine learning model is unable to capture the necessary hardware level and software optimizations on out of distribution dimensions and GPUs. This implies that there are more intricate execution patterns although consistent across the GPUs, but are not easily captured by machine learning models. Thus, we use the machine learning model to make a prediction per tile, and assemble these predictions to determine the latency of the whole operator. This decomposition simplifies the problem into more manageable sub-problems, allowing the machine learning model to be trained on smaller and more straightforward tasks. For example, the typical tile dimensions used by GEMM library ranges from 32 to 256, which is still much smaller than the dimensions of GEMMs in modern DNN workloads. To compute the latency of the entire kernel from the predictions for each tile, we employ the following equations.
@@
Variables $x_{i}$ and $t_{i}$ are the dimensions of the output and the tile, respectively, $N$ is the number of dimensions of the output, and $num_{sm}$ is the number of SMs available in GPU. The tile dimensions are determined by metadata obtained with PyTorch Profiler discussed in Section 6.1. In the next subsection we discuss how we can obtain the $PerTileLatency$.
@@
Next we need to predict GPU’s utilization. We use the observation that a GPU approaches the maximum achievable throughput as the number of available threads in the workload increases (as shown in Figure 5). We formulate the correlation between number of waves and utilization as a machine learning model. This is because, an ML model can capture the non-linear behaviour between kernel latency and GPU characteristics and kernel properties such as operator type, tile dimensions, and arithmetic intensity. We use MLPs as our machine learning model, which is tasked with predicting the coefficients of the following equation:
@@ 4.3. Machine Learning Model to Predict Utilization
We use an MLP with multiple fully-connected layers for our machine learning predictor. We train five distinct MLPs, each tailored for batched matrix multiplication, fully-connected layers, element-wise operators, softmax, and layer normalization.
@@
We use memory size and bandwidth, peak FLOPS, and L2 cache size as GPU features for this modeling. We use these features because: (1) they are publically available even for newer GPUs, (2) bandwidth and peak FLOPs determine the performance bounds of I/O bound and compute bound operators, and (3) on-chip memory size (L2 cache) and off-chip memory size (such as High Bandwidth Memory or DRAM) determine which underlying implementation will be used by the libraries such as cuDNN and CUTLASS to execute the DNN operators. For instance, information for the H100 was made available through NVIDIA’s public documentation after its launch (NVIDIA, 2023a). For the Blackwell architecture, the latest NVIDIA GPU architecture announced (NVIDIA, 2024), details on memory size, bandwidth, and peak FLOPs are already available. Information on the number of SMs and L2 cache size is not availalbe yet, but should become available closer to the release date or shortly after the GPU is released.
@@ 4.4. Support for Operator Fusion
Modern ML frameworks (Chen et al., 2018; PyTorch, a) accelerate the execution of memory-bound kernels by fusing multiple consecutive operators and eliminating the data movement from external memory. NeuSight supports the prediction of the fused operators. When multiple consecutive vector kernels are fused, NeuSight accumulates the FLOPs count of each operator, but discards the memory size required for the intermediate results between kernels. NeuSight uses the updated FLOPs count, memory size of the fused operator, and the metadata including tile size of the first operator to be fused to make predictions with its MLP. For instance, in the GPT-2 architecture, the element-wise addition used for residual connections in each layer can be fused with subsequent layer normalization. NeuSight can predict the latency of this fused operator accumulating the FLOPs count of both operators and discarding the memory size required for the output of the element-wise addition and the input of the layer normalization, and using the predictor for element-wise addition. Another common form of operator fusion is combining the GEMM operator with activation functions. NeuSight supports this fusion using the BMM or fully-connected layer predictor and adjusting the input compute count and memory size.
@@ 5. NeuSight Workflow
Figure 6 shows the overall workflow of NeuSight. NeuSight takes the description of the target deep learning model in PyTorch and extracts the operator/kernel graph using the Torch.fx library (PyTorch, 2023a). It records the metadata of each kernel, including operator type and input/output tensor dimensions. Based on the metadata and tile size, each operator is annotated with predictions made by our kernel-level predictor. The per-GPU latency is computed by accumulating the predicted latency of each kernel executing on a single device, which aligns with device execution where kernels are executed sequentially on the GPU (Zhu et al., 2020).
NeuSight also takes as input the distribution strategy, data parallel width, pipeline parallel depth and schedule, or tensor model parallel width. It is important to highlight that the focus of NeuSight is to effectively predict the performance of DNNs on new or unseen GPUs. Predicting the performance of each kernel is the complex and unsolved aspect tackled by NeuSight, as it involves addressing the non-linearities of GPU software optimizations and hardware execution.
For completeness, we provide a solution that can forecast the latency of distributed execution on a multi-GPU server. The goal of this forecasting is to enable researchers to determine latency of a relatively larger model that often cannot fit on a single device due to memory constraints. This allows NeuSight to provide accurate and reliable predictions for the performance of the compute/memory intensity of the DNN while avoiding the variability introduced by complex multi-server network operations. There are prior works (Wang et al., 2023; Rashidi et al., 2020) that already model complex network topologies, which researchers can leverage with NeuSight.
@@ 5.1. Forecasting for Distributed Execution
Tensor model and Data parallel. NeuSight inserts all-reduce operators into the DNN graph to synchronize the activations (in tensor model parallel) or gradients (in data parallel) across GPUs. The framework currently supports Megatron tensor model strategy (Nvidia, ). The latency of ring-based all-reduce is estimated and aggregated with the compute latency of kernels to determine the end-to-end latency.
@@ 6.1. Implementation Details
Table 3. List of GPUs used to train and test the frameworks.
GPU Year Peak FLOPS Memory Memory # L2 Cache
@@
Hardware. Table 3 shows the GPUs used to evaluate NeuSight. We use 5 GPUs released from 2016 to 2020 for training our MLP predictor. Additionally, we use 2 GPUs released in 2022 and 2023 as out-of-distribution GPUs. For all the final results, A100-40GB variant was used for training but the 100-80GB variant with larger memory and memory bandwidth was used as the testset.
Generating the training dataset. We use the 5 training set GPUs shown in Table 3 to collect data to train our MLPs. Latency is measured by running each operator 25 times and averaging the results. Input tensors are initialized using a normal distribution. All measurements use PyTorch 2.1.0 compiled with CUDA Library 12.1. The collected data covers the following operators:
@@
  • Fully-connected Layer - 32256 data points. Batch size from 1 to 8192, input and output sizes from 1 to 65,536.
@@
We reserve 20% of this sampled dataset for validation. For other operators, such as the embedding layer, which do not contribute significantly to the end-to-end latency (e.g. less than 0.1% for BERT Large), we assume that they are memory-bound operators and estimate latency by dividing the working memory size by the memory bandwidth of the GPU.
Baselines. We compare NeuSight with three baselines: (1) the roofline analysis, most widely used analytical performance estimation method, (2) Habitat, a recent work that uses an MLP to predict GPU performance, and (3) Li et al.  (Li et al., 2023), the linear regression based method to predict the latency of deep learning models. For fair comparisons, we train the baselines with the up-to-date training dataset as NeuSight. For Habitat, we train the MLP following the methodology presented in their paper, but with the larger MLP for higher accuracy (Section 3.2). For element-wise operators, Habitat requires a reference latency of the operator on an existing GPU. We measure the latency of element-wise operators on V100 and use the methodology presented in the paper to predict the latency for other GPUs. For V100, we use the latency measured on P100. For Li et al., following the method described in the paper, we apply linear regression between GPU memory bandwidth and achieved FLOPS performance. We then infer the achieved FLOPS performance of the unseen GPU using its memory bandwidth.
@@
Tile size. We utilize the PyTorch Profiler (PyTorch, 2023b) to obtain the kernel names and the number of thread blocks. For matrix multiplications, we deduce tile sizes from the metadata appended to the kernel names. For other kernels, we use the number of thread blocks to infer tile sizes. We record the kernel name, input dimensions, GPU features, and tile sizes in a database during training. For prediction, we estimate tile sizes by finding the closest match in the database based on the kernel name, input dimensions, and GPU features.
@@
DNN workloads evaluated. Table 4 shows the 6 commonly used model architectures released between 2018 and 2022. For the Switch Transformer, which is a mixture of experts model, a 4-expert configuration is employed, with only one expert active at a time. For each of these models, we measure the inference latency for a binary classification task using BERT, and for a text generation task using GPT-2, GPT-3, OPT, and Switch Transformer. We use the time to generate the first token as the latency metric for text generation tasks. We also report and compare the per-iteration training time, including a single forward and backward pass. Predictions for each model are then compared with the measured latencies to evaluate our framework. Due to the high memory requirement of training, we measure training latency only on GPUs with at least 24 GB of HBM.
6.2. Results

### v2

@@ 3.3. Other Related Works
Performance modeling using analytical frameworks. Several works (Hong, 2009; Qi et al., 2017; Isaev et al., 2023; Hsia et al., 2024) propose analytical models to predict GPU performance. Paleo (Qi et al., 2017), Calculon (Isaev et al., 2023), and MAD-Max (Hsia et al., 2024) decompose latency into computation and communication for distributed training. They estimate latency by measuring effective peak throughput with representative workloads on the target GPU, which requires access to the target GPU for estimation. The primary goal of these works is to estimate performance for distributed execution and employ simpler analytical tools for per-device performance estimation. In contrast, NeuSight focuses on providing a higher-fidelity estimation tool for per-device performance prediction. Nevertheless, NeuSight can be integrated with networking simulators to estimate for distributed settings.
Cycle accurate simulators. Cycle accurate simulators (Bakhoda et al., 2009) require detailed modeling of each GPU component. For every new GPU, the simulator must be updated to support new architectural features, which might not be readily available at launch. Proprietary architectural features can create a gap between real hardware and simulators. Additionally, cycle accurate simulators are not suitable for end-to-end ML application execution due to their long simulation times. While certain works, such as sampled principal kernels (Avalos Baddouh et al., 2021), reduce simulation time, they still take at least hours to days to simulate modern DNN due to the need for modeling every detail of the hardware.
Machine Learning based approaches. Due to the large number of compute lanes in GPUs, they often observe linear performance trends for compute intensive kernels, and hence many prior works leverage regression models to predict the performance of a GPU architecture (Jia et al., 2012; Wu et al., 2015; Li et al., 2023). Some works require extracting low-level hardware features, such as hardware counter values, to make predictions. However, these values are unavailable unless the device is accessible (O’Neal et al., 2017; Punniyamurthy et al., 2017; Wang et al., 2019). Such approaches cannot capture the complex intricacies of hardware features and software optimizations such as tiling for deep learning and do not support cross-generational performance prediction with GPUs due to the lack of hardware. Other works (Yang et al., 2022) propose graph neural networks to predict resource consumption (GPU utilization, memory utilization, and network TX/RX throughput) in distributed deep learning. However, this approach necessitates target GPUs for transfer training.
@@
4. NeuSight Forecasting
NeuSight forecasts the latency of executing a model on multiple GPUs in a server in three steps. The first and main step forecasts the latency of a single kernel that executes on the GPU atomically. The kernel leverages software and hardware optimizations for GPUs, namely tiling, parallel execution across SMs, on-chip caching, and HBM bandwidth. NeuSight uses architectural details such as memory size, memory bandwidth, number of SMs, L2 cache size, and peak FLOPS to predict the kernel’s latency through a machine learning approach. Our kernel estimation is inspired by how deep learning kernels are disassembled into multiple smaller tiles and mapped onto GPU hardware for execution. This approach contrasts with prior works that rely solely on compute or memory utilization to predict latency. Section 4.1 showcases the execution of a few representative DNN kernels on the GPU. NeuSight supports tensor operators such as General Matrix Multiplication (GEMM), element-wise operators such as addition and ReLU, and reduction-based operators such as Softmax and layer normalization. In the second step, the per-operator latency is aggregated to predict the performance of the DNN workload based on its dataflow graph. In the third step, detailed in Section 5.1, the performance of per-GPU execution is combined with predicted network latencies to determine the performance of a DNN executing across multiple GPUs in a server.
@@ 4.1. Kernel Execution on GPUs
Figure 4. Each tile is distributed evenly across SMs and executed concurrently, in multiple number of waves. (NVIDIA, 2023a)
Tiled execution. In this section, we focus on GEMM, a commonly found kernel in DNNs, but NeuSight can predict the latency of various operators to determine the end-to-end performance. GEMM is a core building-block for many deep learning layers, such as fully-connected, convolution, recurrent neural network, or self-attention found in modern large language models. Figure 3 shows a simplified view of a typical dataflow of GEMM on a GPU. Modern GPU libraries execute GEMM by first partitioning the output matrix into multiple identical tiles (Figure 3a). Each tile corresponds to a segment of the output matrix, loads the necessary input operands from input matrices, and computes the output elements. These tiles are mapped to each SM of the GPU and executed concurrently (Figure 3b). Typically the number of tiles which can be executed concurrently is limited by the number of SMs on the GPU, and the entire GPU kernel is executed in multiple waves of tile groups as depicted in Figure 4. Tiling strategy enables scalable execution of GEMM, by decomposing GEMM into multiple smaller workloads, which executes the identical computation on different input and output elements.
Fundamental performance laws of GPUs. Runtime of each kernel is bounded by a few features of a GPU, such as peak memory bandwidth or peak FLOPs. Roofline analysis (Williams et al., 2009) provides a straightforward method to estimate the approximate performance of a GPU kernel, taking into account the arithmetic intensity of the kernel. The intensity is computed by dividing number of flops ($flops_{k}$) by size of memory transactions ($mem_{k}$). The roofline bandwidth can be computed as follows:
(1) $\footnotesize K=\frac{flops_{k}}{mem_{k}};~{}~{}roofline_{BW}=min(K\times memBW_{p},flops_{p})$
Where $flops_{p}$ is the peak FLOPs and $memBW_{p}$ is the peak memory bandwidth of GPU. Roofline bandwidth represents the maximum achievable throughput of a kernel on a GPU.
@@
Latency hiding in GPU. Although the roofline bandwidth represents the theoretical maximum bandwidth, each kernel does not always fully utilize the available compute or memory bandwidth. Table 2 shows the hardware utilization of H100 when executing matrix multiplication, with a sequence length of 512 and a hidden dimension of 64, seen in BERT. The table shows that it is still common in machine learning models to not fully utilize compute or memory bandwidth.
Another performance bound in GPU execution is the use of numerous threads to harness parallelism, achieving significantly higher throughput compared to traditional CPU architectures. GPUs can achieve the aforementioned peak hardware performance ($roofline_{BW}$) only when there are enough available software threads to effectively mitigate latency arising from data dependency or memory stalls by scheduling independent threads. Consequently, GPU performance relies heavily on the level of parallelism in the kernels, which is directly proportional to the number of threads that can be independently scheduled in each SM. Figure 5 illustrates the correlation between the number of waves available in each SM and the achievable throughput. As the number of waves increases, each SM is assigned more waves, resulting in a greater number of threads per SM. This allows for effective hiding of stall latency in each thread, thus leading to higher performance.
Table 2. Measured compute utilization of H100 GPU when executing (512x64) x (64x512) matrix multiplication across various batch sizes.
Batch Size 32 64 128 256 512
Peak FLOPS Utilization 53.2% 70.7% 69.4% 72.3% 86.0%
@@ 4.2. Kernel-wise Prediction
Tile-granularity prediction. For each DNN kernel we break it down into multiple identical tiles, instead of directly predicting the latency of the kernel. This insight is motivated from prior works (Yu et al., 2021), where a machine learning model is unable to capture the necessary hardware level and software optimizations on out of distribution dimensions and GPUs. This implies that there are more intricate execution patterns although consistent across the GPUs, but are not easily captured by machine learning models. Thus, we use the machine learning model to make a prediction per tile, and assemble these predictions to determine the latency of the whole operator. This decomposition simplifies the problem into more manageable sub-problems, allowing the machine learning model to be trained on smaller and more straightforward tasks. For example, the typical tile dimensions used by GEMM library ranges from 32 to 256, which is still much smaller than the dimensions of GEMMs in modern DNN workloads. To compute the latency of the entire kernel from the predictions for each tile, we employ the following:
@@
Variables $x_{i}$ and $t_{i}$ are the dimensions of the output and the tile, respectively, $N$ is the number of dimensions of the output, and $num_{sm}$ is the number of SMs available in GPU. The tile dimensions are determined by metadata obtained with PyTorch Profiler discussed in Section 6.1. In Equation 4, we assume that each SM executes one tile at a time. The latency of the kernel typically scales linearly with the number of waves of tiles, and can be modeled as sequential execution of waves of tiles (NVIDIA, 2024a). As GPUs can execute threads from multiple tiles concurrently, this overlap is accounted for and abstracted in Equation 7, which will be discussed in the following subsection. In the next subsection, we discuss how we can obtain the $PerTileLatency$.
@@
Next we need to predict GPU’s utilization. We use the observation that a GPU approaches the maximum achievable throughput as the number of available threads in the workload increases (as shown in Figure 5). We formulate the correlation between number of waves and utilization as a machine learning model. This is because, an ML model can capture the non-linear behavior between kernel latency and GPU characteristics and kernel properties such as operator type, tile dimensions, and arithmetic intensity. We task MLPs with predicting the coefficients of the following equation:
@@ 4.3. Machine Learning Model to Predict Utilization
We use an MLP with multiple fully-connected layers as our predictor. We train five distinct MLPs, each tailored for batched matrix multiplication, fully-connected layers, element-wise operators, softmax, and layer normalization.
@@
We use memory size and bandwidth, peak FLOPS, and L2 cache size as GPU features for this modeling. We use these features because: (1) they are publicly available even for newer GPUs, (2) bandwidth and peak FLOPs determine the performance bounds of I/O bound and compute bound operators, and (3) on-chip memory size (L2 cache) and off-chip memory size (such as High Bandwidth Memory or DRAM) determine which underlying implementation will be used by the libraries such as cuDNN and CUTLASS to execute the DNN operators. For instance, information for the H100 was made available through NVIDIA’s public documentation after its launch (NVIDIA, 2023c). For the Blackwell architecture, the latest NVIDIA GPU architecture announced (NVIDIA, 2024b), details on memory size, bandwidth, and peak FLOPs are already available. Information on the number of SMs and L2 cache size is unavailable, but based on prior years, should become available closer to the release date or shortly after the GPU is released.
@@ 4.4. Support for Operator Fusion
Modern ML frameworks (Chen et al., 2018; PyTorch, 2024b) accelerate the execution of memory-bound kernels by fusing multiple consecutive operators and eliminating the data movement from external memory. NeuSight supports the prediction of the fused operators. When multiple consecutive vector kernels are fused, NeuSight accumulates the FLOPs count of each operator, but discards the memory size required for the intermediate results between kernels. NeuSight uses the updated FLOPs count, memory size of the fused operator, and the metadata including tile size of the first operator to be fused to make predictions with its MLP. For instance, in the GPT-2 architecture, the element-wise addition used for residual connections in each layer can be fused with subsequent layer normalization. NeuSight can predict the latency of this fused operator accumulating the FLOPs count of both operators and discarding the memory size required for the output of the element-wise addition and the input of the layer normalization, and using the predictor for element-wise addition. Another common form of operator fusion is combining the GEMM operator with activation functions. NeuSight supports this fusion using the BMM or fully-connected layer predictor and adjusting the input compute count and memory size.
@@ 5. NeuSight Workflow
Figure 6 shows the overall workflow of NeuSight. It takes the description of the target deep learning model in PyTorch and extracts the operator/kernel graph using the Torch.fx library (PyTorch, 2023b). It records the metadata of each kernel, including operator type and input/output tensor dimensions. Based on the metadata and tile size, each operator is annotated with predictions made by our kernel-level predictor. The per-GPU latency is computed by accumulating the predicted latency of each kernel executing on a single device, which aligns with device execution where kernels are executed sequentially on the GPU (Zhu et al., 2020).
NeuSight also takes as input the distribution strategy, data parallel width, pipeline parallel depth and schedule, or tensor model parallel width. It is important to highlight that the focus of NeuSight is to effectively predict the performance of DNNs on new or unseen GPUs. Predicting the performance of each kernel is the complex and unsolved aspect tackled by NeuSight, as it involves addressing the non-linearity of GPU software optimizations and hardware execution. For completeness, we provide a solution that can forecast the latency of distributed execution on a multi-GPU server. The goal of this forecasting is to enable researchers to determine latency of a relatively larger model that often cannot fit on a single device due to memory constraints. This allows NeuSight to provide accurate and reliable predictions for the performance of the compute/memory intensity of the DNN while avoiding the variability introduced by complex multi-server network operations. There are prior works (Wang et al., 2023; Rashidi et al., 2020) that already model complex network topologies, which researchers can leverage with NeuSight.
@@ 5.1. Forecasting for Distributed Execution
Tensor model- and Data- parallel. NeuSight inserts all-reduce operators into the DNN graph to synchronize the activations (in tensor model parallel) or gradients (in data parallel) across GPUs. The framework currently supports Megatron tensor model strategy (Smith et al., 2022). The latency of ring-based all-reduce is estimated and aggregated with the compute latency of kernels to determine the end-to-end latency.
@@ 6.1. Implementation Details
Table 4. List of GPUs used to train and test the frameworks. GPUs reserved for the test set are highlighted in grey.
Vendor GPU Year Peak FLOPS Memory Memory # L2 Cache
@@
Hardware. Table 4 shows the GPUs used to evaluate NeuSight. We use 5 Nvidia GPUs and 2 AMD GPUs released from 2016 to 2021 for training our MLP predictor. AMD GPUs have a dedicated datapath for FP32 matrix multiplication, and peak FLOPS for matrix multiplication as shown(AMD, 2022). Additionally, we use 2 Nvidia GPUs released in 2022 and 2023 as out-of-distribution GPUs. For all the final results, A100-40GB variant was used for training but the A100-80GB variant with larger memory and higher memory bandwidth was used as the testset.
Generating the training dataset. We use the 5 training set GPUs shown in Table 4 to collect data to train our MLPs. Latency is measured by running each operator 25 times and averaging the results. Input tensors are initialized using a normal distribution. All measurements were conducted in fp32 format, using PyTorch 2.1.0 compiled with CUDA Library 12.1 for NVIDIA GPUs, and PyTorch 2.4.1 compiled with ROCm 6.1 for AMD GPUs. The collected data covers the following operators:
@@
  • Fully-connected Layer - 32,256 data points. Batch size from 1 to 8192, input and output sizes from 1 to 65,536.
@@
We reserve 20% of this sample dataset for validation. For other operators, such as the embedding layer, which do not contribute significantly to the end-to-end latency (e.g. less than 0.1% for BERT Large), we assume that they are memory-bound operators and estimate latency by dividing the working memory size by the memory bandwidth of the GPU.
@@
Baselines. We compare NeuSight with three baselines: (1) the roofline analysis, most widely used analytical performance estimation method, (2) Habitat (Yu et al., 2021), a recent work that uses an MLP to predict GPU performance, and (3) Li et al.  (Li et al., 2023), the linear regression based method to predict the latency of deep learning models. For fair comparisons, we train the baselines with the up-to-date training dataset as NeuSight. For Habitat, we train the MLP following the methodology presented in their paper, but with the larger MLP for higher accuracy (Section 3.2). For element-wise operators, Habitat requires a reference latency of the operator on an existing GPU. We measure the latency of element-wise operators on V100 and use the methodology presented in the paper to predict the latency for other GPUs. For V100, we use the latency measured on P100. For Li et al., following the method described in the paper, we apply linear regression between GPU memory bandwidth and achieved FLOPS performance. We then infer the achieved FLOPS performance of the unseen GPU using its memory bandwidth.
@@
Tile size. We utilize the PyTorch Profiler (PyTorch, 2023a) to obtain the kernel names and the number of thread blocks. For matrix multiplications, we deduce tile sizes from the metadata appended to the kernel names. For other kernels, we use the number of thread blocks to infer tile sizes. We record the kernel name, input dimensions, GPU features, and tile sizes in a database during training. For prediction, we estimate tile sizes by finding the closest match in the database based on the kernel name, input dimensions, and GPU features.
@@
DNN workloads evaluated. Table 5 shows the 6 commonly used models released between 2018 and 2022. For the Switch Transformer, which is a mixture of experts model, a 4-expert configuration is employed. For each of these models, we measure the inference latency for a binary classification task using BERT, and for a text generation task using GPT-2, GPT-3, OPT, and Switch Transformer. We use the time to generate the first token as the latency metric for text generation tasks. We also report and compare the per-iteration training time, including a single forward and backward pass. Predictions for each model are then compared with the measured latencies to evaluate our framework. Due to the high memory requirement of training, we measure training latency on GPUs with at least 24 GB of HBM.
6.2. Single-Device Execution Results